OpenStudy (anonymous):

Translate this word problem as a system of equations and then solve using substitution Find two numbers whose sum is 58 and whose difference is 16. @InstagramDoll

OpenStudy (mathstudent55):

Let one number be x. Let the other number be y. The sum of the two numbers is x + y, and we know it equals 58, so we have x + y = 58 The difference of the two numbers is x - y, and we are told it's 16, so we have x - y = 16 The system of equations is x + y = 58 x - y = 16 Are you familiar with the substitution method of solving a system of equations?

OpenStudy (anonymous):

wht is the answer

OpenStudy (mathstudent55):

Solve the first equaqtion for x, and substitute into the second equation. x = 58 - y 58 - y - y = 16 -2y = -42 y = 21 Now substitute y = 21 intot he firs equation: x + 21 = 58 x = 37 Answer: x = 37, y = 21